### 15 Mar 2026 v1

Abstract
This article studies phonetic texture in Persian poetry as a literary-historical problem rather than as a by-product of meter or an auxiliary feature for classification. The corpus comprises 1,116,306 mesras from 31,988 poems and 83 poets drawn from a larger archive after restricting inference to five canonically named classical meters in the retained analytical subset and to well-supported poet-meter cells. Each mesra is represented as a grapheme-to-phoneme symbol stream and summarized through six measures: hardness, sonority, sibilance, vowel ratio, phoneme entropy, and consonant-cluster ratio.
The argument combines corpus description, confound-aware modeling, and literary interpretation. Poet contrasts are estimated under controls for meter, poetic form, and line-length opportunity; century trends are interpreted as historical redistribution rather than rupture; and meter is treated as a prosodic institution that shapes which phonetic differences can stabilize within a line. Meter and form explain substantial variance, yet they do not absorb poet-level differentiation. Persian poetic sound is therefore best understood as conditioned variation within shared formal regimes rather than as either pure authorial essence or mere metrical residue.
The literary implications emerge most clearly in the stylistic map and case studies. The phonetic space separates high-sonority lyrical profiles, hardness-led rhetorical or epic profiles, sibilant mystical and courtly contours, and high-entropy experimental textures. Century analysis links early firmness to epic and courtly baselines, later sonority to the prominence of ghazal and mystical lyric, and sixteenth- to nineteenth-century reconfiguration to Indo-Persian complexity, revivalist rhetoric, devotional address, and changing performance contexts. Case studies of Safi Alishah, Shah Nematollah Vali, Bidel Dehlavi, Ashofteh Shirazi, Fakhr Al-Din Asad Gorgani, Ferdowsi, Saadi, Khaqani, Jahan Malek Khatun, Bolandeghbal, Hafez, Shahriar, and Khayyam show that familiar literary labels such as firmness, luminosity, difficulty, and fluency correspond to different phonetic mechanisms in different historical settings.
The article makes four contributions: it establishes a corpus-scale phonetic framework for Persian poetry; models poet differentiation under explicit formal controls while keeping meter legible as classical aruz; develops a multidimensional account of stylistic clustering and historical redistribution rather than a single hard-soft continuum; and demonstrates how computational phonetics can sharpen literary-historical interpretation without displacing it.
keywords:
digital humanities, Persian poetry, computational poetics, historical poetics, corpus stylistics
1 Introduction
Sound has long been one of the most durable critical languages of Persian literary culture. Epic verse is heard as firm and declarative; courtly qasida as sharpened and brilliant; ghazal as open, memorable, and singable; mystical poetry as fluid, repetitive, and incantatory; later Indo-Persian writing as dense and internally restless; modern lyric as more conversationally spacious. These descriptors are not incidental ornaments. They are shorthand for how poetic language distributes consonantal pressure, vocalic openness, strident contour, and rhythmic closure. Yet they are difficult to compare beyond a limited set of canonical poems. Close reading can hear the sonic pressure of a hemistich, but not easily decide whether that pressure belongs to one line, one meter, one poet, or a broader historical formation.
Persian makes the problem especially demanding. The script ordinarily omits short vowels, so purely graphemic comparison is insufficient for a study of poetic sound. At the same time, meter cannot be treated as a weak background variable. In Persian verse, aruz is a constitutive institution: it shapes diction, line architecture, and the range of sonic expectations available to poets. Here meter means a named bahr: a quantitative pattern of long (-) and short ($\cup$) syllables distributed across the mesra and realized through historically accepted sub-patterns and zehafat. Poetic form is a different category, referring instead to poem-level organizations such as ghazal, masnavi, qasida, or rubai. A serious corpus study of sound must therefore do two things at once. It must reconstruct a stable phonetic surface from the written text, and it must distinguish poet-level variation from the variation made available by meter, form, and line length. That double requirement turns phonetic analysis into both a methodological problem and a humanities problem.
Recent Persian digital humanities has supplied much of the necessary infrastructure. Large literary corpora, poet-level stylometry, meter recognition, and school detection have made comparative work increasingly feasible. What remains underdeveloped is phonetic texture as an interpretive object in its own right. Sound features often appear inside hybrid authorship or classification models, but they are rarely treated as evidence for literary-historical comparison, and even more rarely under explicit control for the formal structures that organize Persian verse. This gap matters because some of the oldest judgments in Persian criticism are precisely sonic judgments. If computation cannot address them responsibly, it leaves a central dimension of literary form understudied.
This article asks four linked questions. First, do poet-level phonetic differences remain measurable after controlling for meter, poetic form, and line-length opportunity? Second, how are those differences organized in a multidimensional stylistic field rather than along a single hard-soft axis? Third, what century-by-century redistributions become visible when phonetic measures are read against Persian literary history? Fourth, how can those quantitative profiles be translated into historically credible readings of individual poets without collapsing literary value into sound alone?
The study answers those questions through a deliberately disciplined design: corpus restriction for comparability, symbolic phonetic representation at mesra scale, confound-aware statistical models, century interpretation, and case studies that return the measurements to literary history. The active inferential cohort comprises 1,116,306 mesras from 31,988 poems and 83 poets. Ten of the main case-study poets fall squarely inside that balanced cohort; Ferdowsi is added as a high-support epic reference; Shahriar and Khayyam are retained as descriptive extensions whose historical importance exceeds their fit with the strongest inferential frame. This architecture lets the article compare poets rigorously while remaining explicit about evidential boundaries. The article makes four contributions:
  1. 1.
    It develops a corpus-scale phonetic analysis of Persian poetry at mesra resolution, using a grapheme-to-phoneme representation suitable for large literary comparison.
  2. 2.
    It estimates poet differences under explicit controls for meter, poetic form, and line-length opportunity, while translating internal meter codes into historically legible classical aruz names.
  3. 3.
    It models Persian poetic sound as a multidimensional stylistic field organized around openness, firmness, contour, and entropy, and reads that field across centuries and across named metrical environments rather than as a single hierarchy of hardness or softness.
  4. 4.
    It demonstrates, through historically grounded case studies, how phonetic metrics can sharpen interpretations of epic, ghazal, mystical verse, courtly rhetoric, revivalist poetics, and modern lyric without reducing poetry to phonetics alone.
2 Related Work
Research on Persian poetry in digital humanities is now rich enough that the main question is no longer whether large-scale comparison is possible, but what kind of literary object each computational design actually makes visible. The present study belongs at the intersection of corpus infrastructure, stylometry and prosody, phonetic representation, and sound-centered digital humanities, but it addresses a gap that those strands have not yet resolved together.
2.1 Persian Corpora and Literary Modeling
Corpus construction has been the enabling condition of recent Persian DH. Early work by Asgari and Chappelier showed that Persian poetic corpora could sustain large-scale analysis of literary tendencies rather than only isolated text-processing tasks [1]. That infrastructural turn was extended decisively by Raji et al., whose large corpus and metadata architecture made poet-, genre-, and period-level comparison practically viable [2]. More recent systems such as PARSI and NAZM have moved from corpus assembly to literary profiling, showing that Persian poets and traditions can be compared through stylometric, semantic, and metrical features [3, 4]. Work on thematic similarity and era-sensitive classification likewise demonstrates that Persian poetic corpora can sustain temporal and stylistic discrimination at poem scale [5, 6]. This literature establishes the feasibility of poet-level modeling, but its primary objects are usually authorship, chronology, theme, or broad style rather than phonetic surface as a humanities object in its own right.
2.2 Prosody, Stylometry, and Authorship
The closest methodological neighbors to this article come from stylometry, authorship studies, computational prosody, and phonetic representation. Quantitative work on attribution, such as Ishikawa’s study of the Attar corpus, demonstrates how statistical evidence can enter philological debate without claiming to replace it [7]. Prosodic studies likewise show that Persian aruz is computationally tractable, as in Shahrestani and Chehreghani’s recognition work [8]. At the level of phonetic reconstruction, Qharabagh et al. demonstrate that Persian grapheme-to-phoneme conversion is itself an active computational problem rather than a transparent preprocessing step [9]. Parallel work in Arabic and multilingual meter modeling reinforces the same point: formal structure can be learned computationally and matters deeply for literary analysis [10, 11].
What these studies do not yet offer is a phonetic stylistics in which meter is both controlled and interpreted. A meter recognizer identifies formal pattern; a G2P benchmark secures representational plausibility; an authorship model may use sound features instrumentally; none necessarily explains how poets working inside a shared meter redistribute hardness, sonority, or sibilant contour for different literary ends. For Persian, where metrical repertoires strongly condition lexical and rhythmic opportunity, that distinction is crucial.
2.3 Sound-Centered Digital Humanities
Outside Persian studies, sound-centered digital humanities has provided both conceptual orientation and methodological precedent. Clement’s call to “sound for meaning” argued that aural patterning belongs to interpretation rather than to a merely decorative supplement [12]. Poemage turned that principle into an exploratory visual environment for sonic structure within poems [13]. More recently, Meng et al. showed that phonetic vectors can distinguish major Chinese poets under strong formal constraint, demonstrating that sound can support historically specific comparison instead of only generic euphony claims [14].
These studies are important precisely because they show both the promise and the limit of computational sound analysis. They confirm that sound is measurable and interpretable, but they often privilege local visualization, poet-pair contrast, or predictive discrimination over a sustained account of formal confounds. Persian requires a stricter design because its script under-specifies vowels and its major poetic traditions are inseparable from prosodic institutions.
2.4 Gap Addressed Here
The gap addressed by this article is therefore specific. Persian DH already has strong work on corpus infrastructure, stylometry, prosody, phonetic engineering, generative modeling, and broader AI-oriented literary exploration [15, 16, 17, 18, 9]. Sound-centered DH already shows how phonetic patterning can matter for literary interpretation. What is missing is a Persian framework that treats phonetic texture itself as the object of explanation, while controlling explicitly for meter, form, and line-length opportunity and then returning the results to literary history.
This study addresses that gap by shifting the inferential target. Sound is not used primarily as an auxiliary predictor for classification. Instead, it is modeled as a multidimensional field in which poets, meters, and centuries occupy historically intelligible positions. Its novelty lies not in rejecting prior Persian DH, but in assembling its strongest strands into a single argument: phonetic representation, formal control, meter interpretation, stylistic clustering, and literary-historical reading belong in the same analytical frame if we want to hear the corpus historically rather than merely process it computationally.
3 Data and Preprocessing
The article begins from a large mesra-level archive, but it does not treat the archive as one homogeneous evidential field. In Persian poetry, phonetic comparison is inseparable from meter, poetic form, and line length; preprocessing is therefore part of the scholarly design, not merely a technical prelude. The aim is to define a domain in which poet-level comparison remains historically legible and formally controlled.
3.1 Corpus Scope and Inferential Domain
The broad corpus contains 2,892,576 mesras extracted from poem-level source records. The controlled analytical population used for the main models is smaller: 1,116,306 mesras from 31,988 poems and 83 poets. This restriction is deliberate. Rows without usable meter or form metadata cannot sustain confound-aware comparison, and very sparse poet-meter cells do not support credible within-meter interpretation.
The inferential cohort is further limited to verse in five canonically named classical meters of Persian aruz: mutaqarib, hazaj, ramal, mujtass, and muzari’. In Persian prosody, a bahr is a line-level quantitative pattern of long (-) and short ($\cup$) syllables distributed across the mesra; individual realizations admit historically recognized sub-patterns and zehafat rather than free variation. [^1] For the canonical nomenclature of Persian quantitative prosody, including ramal, hazaj, rajaz, mutaqarib, muzari’, mujtass, khafif, munsarih, and muqtadab, see [19, 20, 21]. Spellings such as mutagharib for mutaqarib, modar for muzari’, mujtath for mujtass, and maqtadab for muqtadab reflect transliteration choice rather than distinct meters. In this article the subsequent five-way grouping is a post-hoc analytic convenience: the canonical literary category is the named bahr, while corpus codes such as M01–M05 identify specific retained sub-patterns. These are not arbitrary computational bins. They are historically documented meters that dominate the metrically annotated archive and form the part of the corpus in which controlled comparison remains statistically credible and literarily legible.
Table 1: Construction of the inferential domain from the broad corpus. Main-text models are estimated only on the final cohort.
Stage Mesras
Broad mesra layer after structured extraction 2,892,576
Rows with usable meter/form metadata inside the five retained classical meters 1,439,014
Final inferential cohort after imposing a minimum of 2,000 mesras per poet-meter cell 1,116,306
The final cohort retains 38.6% of broad-corpus mesras and 83 of 216 poets. Read from the metrically annotated subset rather than the full archive, the five retained meters account for 77.6% of all mesras with usable meter/form metadata, and after the support threshold they constitute the entire inferential cohort. Their shares inside the final cohort are 30.3% for mutaqarib, 21.1% for hazaj, 17.8% for ramal, 17.3% for mujtass, and 13.5% for muzari’. The gain is not statistical tidiness alone. It is literary discipline: poets are compared inside shared formal conditions, and later interpretation is not driven by marginal meter cells, unbalanced century tails, or incomplete metadata.
3.2 Normalization, Alignment, and Identity Control
Preparation of the corpus is intentionally conservative. Orthographic normalization harmonizes a limited set of high-impact variants, especially Persian and Arabic forms of ya and kaf, tatweel, and spacing inconsistencies, while preserving line structure. Mesra boundaries are accepted only where structurally explicit, and poet identities are canonicalized across variant spellings so that author labels and century metadata do not fragment. These operations matter less as file handling than as literary safeguards: unstable line boundaries or split author identities would distort exactly the comparisons on which later historical interpretation depends.
Two further restrictions define the main analytical domain. First, the study remains inside the five most frequent named classical meters because these meters provide the support needed for controlled estimation and meaningful within-meter reading. This is a statistical restriction rather than a canonical literary taxonomy: the analytical grouping is post hoc, whereas the historically documented category is the named bahr and, where necessary, its specific sub-pattern. Meter and poetic form are also kept analytically separate. Meter specifies the quantitative line pattern, whereas form refers to poem-level organizations such as ghazal, masnavi, qasida, and rubai; the former governs rhythmic slotting within the mesra, while the latter shapes generic expectation, rhyme architecture, and discourse scale. Second, line-length descriptors are retained so that longer lines do not mechanically generate apparent stylistic differences. These operations do not make the corpus neutral. They make it comparable.
3.3 Corpus Structure
Figure 1 now summarizes the analytical population through two coordinated views. A figure header reports the scale of the balanced cohort directly, avoiding the distracting indexed-count panel used in earlier drafts. Panel A turns to prosody by showing that the comparison set is concentrated in the five retained classical meters. Panel B then shows the historical coverage of the metadata-linked subset, which is strongest from the tenth through seventeenth centuries and preserves a smaller but important nineteenth-century tail.
Figure 1: Structure of the analytical corpus. The header reports the scale of the balanced cohort directly: 1,116,306 mesras, 31,988 poems, and 83 poets across the five retained classical meters. In mesra terms these meters account for 77.6% of rows with usable meter/form metadata and 100% of the final inferential cohort. Panel A shows poem counts and shares across mutaqarib, hazaj, ramal, mujtass, and muzari’, identifying the prosodic field within which controlled comparison is possible. Panel B plots poems per century for the metadata-linked subset, showing the historical span on which the century analysis rests.
Taken together, these distributions clarify the interpretive stance of the article. The models are not estimating “Persian poetry” in the abstract. They are estimating phonetic differentiation inside a historically broad but formally disciplined subset dominated by a small number of aruz repertoires. That is a narrower claim than canon-wide representation, but a much stronger claim than unrestricted corpus description.
3.4 Interpretive Consequence
The preprocessing argument is therefore straightforward. A large corpus becomes a persuasive humanities object only when the boundaries of comparison are as explicit as the scale of the archive. By restricting the inferential domain, the article gives up some breadth in order to gain historical legibility: epic diction is compared with lyric diction inside identifiable meters, later poets are not allowed to dominate thinly supported cells, and century claims remain tethered to known coverage. That trade-off is essential to everything that follows.
4 Phonetic Representation
Persian orthography ordinarily omits short vowels, so graphemic comparison alone is inadequate for a study of poetic sound. The corpus is therefore represented through a grapheme-to-phoneme layer that approximates pronunciation in a structured symbolic form. This representation remains symbolic rather than acoustic: it does not model recitational performance, dialectal variation, or historical vocal delivery. What it provides is a consistent phonetic surface on which large-scale comparison becomes possible.
4.1 Symbol Streams and Boundary Rules
Each mesra is converted into an ordered sequence of atomic symbols after minimal cleanup. Word boundaries are preserved, mesra boundaries are split only on explicit structural delimiters, and in-token phonological symbols are not mistaken for separators. This rule is essential in Persian G2P material, where some characters can function as phonological units in one context and as boundary markers in another. The representation is designed for transparency rather than compression: later claims about elevated sibilance or reduced sonority remain traceable to readable segmental distributions.
4.2 Feature Mapping
Each symbol is associated with a compact feature profile describing broad articulatory or acoustic behavior: segment type, voicing, stridency, and the scalar values used here for hardness and sonority. The sonority scale follows a standard hierarchy in which stops receive the lowest value (1.0), fricatives 2.0, nasals 3.0, liquids approximately 3.5–3.7, glides approximately 4.1–4.2, and vowels 5.0. The hardness scale is intentionally asymmetric: vowels sit at 0.5, liquids and nasals mostly between 1.2 and 2.0, voiced fricatives near 2.5–2.8, dental stops near 3.0–4.0, and dorsal or uvular obstruents at the high end (up to 4.9). These weights are heuristic rather than acoustic absolutes. Their purpose is not to recover historical sound perfectly, but to provide a stable comparative scheme across a very large corpus.
The weighting scheme matters interpretively. A poet can become “harder” either by accumulating more obstruent pressure overall or by favoring specifically dorsal and uvular segments; conversely, sonority can rise through vowels and glides without eliminating all consonantal edge. The paper therefore treats the representation as an interpretable feature map rather than as an opaque embedding.
4.3 Mesra as Analytical Unit
The mesra is the primary unit of analysis because it balances local phonetic texture and poetic form. Word-level analysis is often too narrow for stylistic inference, while poem-level aggregation can conceal the internal distribution of sound. Mesra-level representation preserves line-scale pressure and still allows later aggregation by bayt, poem, poet, or century.
Two structural descriptors are carried forward with the representation: total symbol count and token count. These are not stylistic outcomes. They are controls that help distinguish phonetic patterning from line-length opportunity.
4.4 Representational Limit
The representational claim of the article is precise. Hardness, sonority, sibilance, and related measures describe encoded phonetic texture in the corpus representation, not direct recordings of historical recitation. This limit is important, but it does not weaken the study. It clarifies the inferential object. Once line boundaries, named meters, poet identities, and phonetic symbols are stabilized, corpus statistics can be returned to literary history without pretending to measure performance itself.
5 Metrics and Statistical Models
The analysis models Persian poetic sound through six complementary measures defined over the mesra-level symbol stream. The point is not to collapse style into a single acoustic index, but to preserve several dimensions that criticism often gathers under broader labels such as firmness, fluency, brilliance, density, or softness.
5.1 Phonetic Metrics
Three measures serve as the primary modeled outcomes. Hardness is the mean symbol-level hardness score across a mesra,
$\text{Hardness}=\frac{1}{n}\sum_{j=1}^{n}h_{j},$
and functions here as a proxy for consonantal firmness or articulatory closure. Sonority is the mean symbol-level sonority score,
$\text{Sonority}=\frac{1}{n}\sum_{j=1}^{n}s_{j},$
and captures relative openness, vocalic breadth, or ease of phonetic flow. Sibilance is the proportion of strident consonants among consonantal symbols,
$\text{Sibilance}=\frac{\#\text{strident consonants}}{\#\text{consonants}},$
and isolates hiss-like contour independently of overall vowel load. In literary terms, these three dimensions correspond roughly to the resources often heard as epic or didactic firmness, lyric openness, and rhetorical or mystical sharpness.
Three additional measures are used mainly for descriptive profiling. Vowel ratio records vocalic occupancy in the line, consonant-cluster ratio measures local compaction through adjacent consonant sequences, and phoneme entropy measures the internal diversity of symbol usage within the mesra:
$\text{Entropy}=-\sum_{k}p_{k}\log_{2}p_{k}.$
These secondary metrics matter because poets who look similar on one primary dimension can diverge strongly on the others: two poets may both sound “open,” for example, while differing sharply in cluster pressure or entropy.
For visualization, poet-level means are normalized to a shared 0–1 range within each metric so that fingerprint plots compare relative position rather than raw scale. Statistical models, however, are fit on the unnormalized metric values.
5.2 Primary Controlled Specification
The inferential core estimates poet contrasts under explicit formal control. For each primary outcome $Y\in\{\text{hardness},\text{sonority},\text{sibilance}\}$, the main model is
$Y_{i}=\alpha+C(\text{poet}_{i})+C(\text{meter}_{i})+C(\text{form}_{i})+\lambda_{1}\,\text{symbols}_{i}+\lambda_{2}\,\text{tokens}_{i}+\varepsilon_{i}.$
This specification addresses the central confound of quantitative poetics: a poet may appear distinctive simply because that poet’s corpus concentrates in one meter, one form, or one line-length regime. Observations remain mesra-level, while uncertainty is clustered at the poem level to account for within-poem dependence.
Meter and form enter separately because they operate at different formal scales. Meter specifies the line-level quantitative pattern of long and short syllables and its permissible sub-patterns, whereas form distinguishes poem-level organizations such as ghazal, masnavi, qasida, and rubai. The former conditions rhythmic slotting and phonotactic opportunity inside the mesra; the latter shapes discourse length, rhyme architecture, and generic expectation across the poem.
5.3 Modeling Strategy
Three additional modeling choices organize the argument. First, nested hardness models estimate the incremental contribution of line length, meter and form, and poet identity. This distinguishes poet signal from formal structure instead of treating all explained variation as interchangeable. Second, century trends are summarized through reduced-control regressions and poem-bootstrap descriptive estimates, because the historical question concerns long-range redistribution rather than one-to-one poet comparison. Third, within-meter models are fit for the five dominant named classical meters so that poet differences can be read inside a shared prosodic envelope: mutaqarib, hazaj, ramal, mujtass, and muzari’. Those meters are retained not because they exhaust Persian aruz, but because they constitute the historically documented and statistically dominant part of the metrically annotated corpus.
5.4 Interpretive Hierarchy
The paper maintains a clear evidential hierarchy. Hardness, sonority, and sibilance support the strongest inferential claims because they are modeled directly. Vowel ratio, entropy, and cluster ratio mainly clarify how poet profiles are assembled. Principal-component projection and fingerprint plots are therefore descriptive aids, not independent inferential results. They matter because they make the geometry of style readable, but they are always interpreted against the controlled models, the meter profiles, and the historical argument.
That hierarchy is essential for responsible reading. Statistical significance alone is not treated as a literary claim. When the paper moves from coefficients to interpretation, it does so through effect magnitude, cross-metric coherence, meter context, and historical plausibility. The method is designed to discipline literary judgment, not to mechanize it.
6 Results
The results move from controlled differentiation to historical interpretation. The first layer establishes that the core phonetic measures are well behaved and that poet effects persist after formal control. The second reads century trends against the actual temporal coverage of the corpus: the tenth through seventeenth centuries, with a smaller nineteenth-century return. The third places poets in a multidimensional stylistic space, and the fourth returns meter from control variable to prosodic institution.
6.1 Global Patterns and Controlled Contrasts
Figure 2 shows that hardness, sonority, and sibilance are all sufficiently dispersed to support comparison. Hardness and sonority occupy dense central ranges, whereas sibilance is more selective and right-skewed. In literary terms, this means that strong strident contour is not the default condition of Persian verse; it is a marked resource used more intensively by particular poets and traditions.
Figure 2: Distributions of the three primary phonetic outcomes in the controlled analytical population. Hardness and sonority cluster around stable central tendencies, whereas sibilance has a longer upper tail, indicating that strongly strident texture is a specialized rather than ordinary condition.
The next question is whether formal structure exhausts that variation. Figure 3 shows that line length, meter, and form all contribute meaningful explanatory power in the nested hardness models. They are therefore indispensable controls, not optional refinements. Yet poet identity still adds explanatory value after those layers are imposed. The result is neither romantic authorial essence nor formal determinism. It is constrained differentiation within shared institutions.
Figure 3: Incremental variance explained by successive control layers in the hardness models. Meter and form account for substantial structure, but poet identity still contributes additional signal once those layers are controlled.
Figure 4 makes that residual layer concrete. Shah Nematollah Vali occupies the strongest negative hardness pole in the balanced cohort ($\beta=-0.062$), while Saadi stands among the clearest positive poles ($\beta=0.034$); Asadi Tusi, Kamal al-Din Ismail, and Jamal al-Din Abd al-Razzaq Isfahani reinforce the same firm rhetorical side of the field. These are moderate effects, and their moderate scale is part of their significance. The models do not reveal acoustically isolated worlds. They reveal recurrent stylistic tendencies inside a common prosodic ecology.
Figure 4: Selected controlled hardness contrasts with confidence intervals. The coefficients remain moderate in magnitude, indicating that Persian poetic style is neither metrically exhausted nor acoustically discontinuous across poets.
6.2 Century-by-Century Historical Redistribution
Figure 5 reads historical change as redistribution rather than rupture. Hardness is highest in the tenth-century subset (mean 1.612) and remains relatively elevated in the seventeenth century (1.604), while the twelfth through sixteenth centuries occupy a slightly softer middle band. Sonority rises from the thirteenth to the fifteenth century, peaks in the fifteenth-century subset (3.235), dips again in the seventeenth century, and rises descriptively in the nineteenth-century tail. Sibilance never becomes the dominant dimension of the corpus as a whole, but it is visibly stronger in the fourteenth and nineteenth centuries than in the early baseline. Entropy rises most clearly from the fourteenth century onward and reaches its maximum in the seventeenth century (3.862).
Figure 5: Century trajectories for hardness, sonority, sibilance, and entropy with poem-bootstrap uncertainty bands. The curves show redistribution across the attested centuries of the corpus rather than a single linear shift from “hard” to “soft.”
These trajectories are historically intelligible only when read century by century. The tenth century combines the highest hardness with relatively low entropy, a profile consistent with early epic and courtly consolidation. That does not mean the early field is uniformly severe: already in the eleventh and twelfth centuries the archive includes romance and panegyric corpora whose sonic organization is more varied than a single heroic baseline would imply. Fakhr Al-Din Asad Gorgani is instructive here, because his hazaj-based romance diction is markedly sonorous inside an otherwise firmer early field.
The thirteenth through fifteenth centuries register the broad centrality of ghazal, mystical lyric, and courtly address. Sonority rises in those centuries, but the rise should not be romanticized as a simple move from hardness to softness. Saadi, Hafez, Jahan Malek Khatun, and Shah Nematollah Vali all participate in lyric or quasi-lyric traditions, yet they reach lyric audibility through different phonetic routes: Saadi through disciplined closure, Hafez through sonority plus contour, Jahan Malek Khatun through bright vocalic spread, and Shah Nematollah Vali through devotional fluency within ramal. What increases historically is not softness as such, but the prestige of line architectures that can sustain openness, chantability, and vocal extension.
The sixteenth and seventeenth centuries register a different transformation. Here entropy and mixed pressure become more salient than raw hardness alone. This pattern suits later Persianate and Indo-Persian poetics, in which density often arises from rapid conceptual turns, lexical heterogeneity, and local internal modulation. Bidel Dehlavi is the clearest poet-level embodiment of that field. His distinctiveness lies less in extreme hardness than in symbolic variety, cluster support, and acoustic swerving within the line. The seventeenth-century rebound in hardness should therefore not be misread as a return to the tenth-century epic baseline; it is a later-classical recombination of pressure with much greater internal diversity.
The nineteenth century appears in the corpus as a smaller but revealing tail. Its higher sibilance and renewed sonority fit a period in which revivalist rhetoric, didactic-public address, devotional circulation, and expanding print culture redistribute older resources rather than simply repeating them. Safi Alishah, Ashofteh Shirazi, and Bolandeghbal exemplify three distinct nineteenth-century solutions: exhortative mystical contour, calibrated bazgasht modulation, and Sufi-didactic public bite. The century curves therefore track changes in literary ecology, performance context, and poetic register, not a teleological sonic destiny.
6.3 Phonetic Stylistic Space and Cluster Logic
Figure 6 projects poet-level means for the six phonetic measures into a low-dimensional stylistic space. PC1 is dominated by vowel ratio and sonority against consonant-cluster pressure, so it is best read as an openness axis. PC2 is dominated by hardness and sibilance, and is therefore better understood as articulated pressure and contour than as hardness alone. The projection is descriptive, but it clarifies the geometry within which the case studies become legible.
Figure 6: Phonetic stylistic space of Persian poets. Grey points show the wider field; colored points mark the focal case studies and historical anchors. The cluster legend is placed outside the plotting area so the labeled cases remain readable.
Four regions matter most. The high-sonority lyrical cluster includes Shah Nematollah Vali, Fakhr Al-Din Asad Gorgani, Jahan Malek Khatun, and descriptively Shahriar; Hafez sits near this zone but rises toward greater contour. The high-hardness rhetorical cluster brings together Ferdowsi, Saadi, and descriptively Khayyam, where authority is carried by closure and compaction more than by bright fricative color. A sibilant rhetorical-mystical region contains Safi Alishah, Khaqani, and Bolandeghbal, but the shared contour masks very different literary functions: exhortative Sufi address, courtly brilliance, and didactic-public insistence. Finally, the high-entropy zone centered on Bidel and Ashofteh marks styles whose distinctiveness lies in internal heterogeneity rather than in any single dominant register.
Table 2: Phonetic archetypes in the Persian stylistic space. Archetypes are descriptive regions rather than discrete school labels, and they summarize the stylistic map discussed in the main text rather than imposing a rigid taxonomy on poets or periods.
Archetype Metric signature Representative poets Literary interpretation
High-sonority lyric High sonority and vowel ratio, low compaction, usually low-to-moderate hardness Shah Nematollah Vali, Fakhr Al-Din Asad Gorgani, Jahan Malek Khatun, Rumi, Shahriar (descriptive) Open vocalic texture, sustained lyric or contemplative cadence, and reduced consonantal congestion without acoustic blandness.
High-hardness rhetorical High hardness with cluster support and restrained sibilance Saadi, Ferdowsi, Asadi Tusi, Khayyam (descriptive) Aphoristic, epic, or didactic firmness in which authority comes from consonantal closure more than from fricative brightness.
Sibilant mystical / rhetorical Elevated sibilance with moderate hardness and strong local articulation Safi Alishah, Khaqani, Hafez, Bolandeghbal Courtly, devotional, or lyric sharpness produced by contour and strident edge rather than by maximal pressure.
High-entropy complexity High entropy with mixed hardness and sonority, often with cluster support Bidel Dehlavi, Ashofteh Shirazi, Saeb, Juya Tabrizi Later Persianate or return-style texture in which symbolic variety and local contrast become the primary vehicles of distinctiveness.
Table 2 summarizes these zones as descriptive archetypes. The point is comparative rather than taxonomic. The table shows that lyric openness, rhetorical force, sibilant brightness, and experimental density recur across traditions, but they do so through historically distinct combinations of the six metrics.
6.4 Meters as Prosodic Institutions
Meter is not only a control variable in this article. It is one of the principal institutions through which sonic expectation is organized. Table 3 translates the internal meter codes into named classical aruz meters and the exact retained sub-patterns, and reports poem-level phonetic means for the five dominant meters in the balanced subset. [^2] In the analytic coding used here, M01–M05 correspond to the sub-patterns mutaqarib-i musamman-i mahzuf, hazaj-i musaddas-i mahzuf, ramal-i musamman-i mahzuf, mujtass-i musamman-i makhbun-i mahzuf, and muzari’-i musamman-i akhrab-i makfuf-i mahzuf. Following [19, 20, 21], these are best understood as specific sub-patterns or permissible variants within named bahrs, not as members of an abstract literary “family.” The five meters were selected because they dominate the metrically annotated archive and therefore support both statistical generalization and historically intelligible interpretation: together they account for 77.6% of mesras with usable meter/form metadata and 100% of the final support-filtered cohort.
Table 3: Most common prosodic meters in the balanced analytical cohort, with poem-level phonetic means. Together the five retained meters account for 77.6% of mesras with usable meter/form metadata and 100% of the final controlled cohort; within that cohort their mesra shares are 30.3% (mutaqarib), 21.1% (hazaj), 17.8% (ramal), 17.3% (mujtass), and 13.5% (muzari’). Internal corpus labels are mapped to standard Persian names for the relevant bahrs and to the exact retained sub-patterns so that the computational codes remain readable in literary-historical terms. Metrics are averaged within poems before meter-level aggregation so that long works do not dominate the comparison.
Code / meter Scansion Retained sub-pattern Poems Mesras Hard. Son. Sib. Ent.
M01 / Mutaqarib fa’ulun fa’ulun fa’ulun fa’ul mutaqarib-i musamman-i mahzuf 3,619 338,250 1.608 3.193 0.179 3.707
M02 / Hazaj mafa’ilun mafa’ilun fa’ulun hazaj-i musaddas-i mahzuf 4,749 236,094 1.582 3.217 0.183 3.715
M03 / Ramal fa’ilatun fa’ilatun fa’ilatun fa’ilun ramal-i musamman-i mahzuf 9,690 198,726 1.579 3.216 0.176 3.873
M04 / Mujtass mafa’ilun fa’ilatun mafa’ilun fa’ilun mujtass-i musamman-i makhbun-i mahzuf 7,334 193,074 1.601 3.247 0.187 3.817
M05 / Muzari’ maf’ulu fa’ilatu mafa’ilu fa’ilun muzari’-i musamman-i akhrab-i makfuf-i mahzuf 6,603 150,652 1.603 3.221 0.186 3.818
Classical prosody and literary history make the same selection intelligible. Mutaqarib is the canonical epic meter of the Shahnameh tradition and remains strongly associated with heroic or declarative narrative; hazaj is prominent in romance, chantable lyric extension, and later devotional reuse; ramal is central to ghazal and mystical lyric from the high-classical period onward; mujtass regularly supports compact ghazal brightness; and muzari’ is especially visible in rhetorically wrought late-classical, Indo-Persian, and revivalist verse. [^3] For the canonical metrical inventory and the status of these forms as named Persian bahrs, see [19, 20, 21]. The genre and period associations emphasized here synthesize those standard descriptions with the distributions observed in this corpus. The article therefore uses historically documented meters rather than approximate metrical classes, while still recognizing that the five-way grouping is an analytic restriction imposed for inferential stability.
The meter profiles are historically legible and phonetically interpretable. The retained mutaqarib sub-pattern, mutaqarib-i musamman-i mahzuf, is the hardest and least sonorous meter in the cohort, which helps explain its long association with epic narrative and declarative force. Hazaj-i musaddas-i mahzuf is softer and more sonorous, fitting its roles in romance, chantable lyric extension, and some devotional repertories. Ramal-i musamman-i mahzuf is the least sibilant and the most entropic retained meter, making it especially hospitable to reflective, mystical, and internally modulated diction. Mujtass-i musamman-i makhbun-i mahzuf combines the highest sonority with the highest sibilance, which helps explain its brightness in ghazal performance. Muzari’-i musamman-i akhrab-i makfuf-i mahzuf retains relatively high hardness while also supporting elevated contour and diversity, making it especially suitable for rhetorically charged late-classical texture. These phonetic tendencies should not be reified as metrical essences, but they are consistent with a simple structural intuition: tighter cadential closure and greater consonantal compression favor hardness, whereas more open vocalic sequencing and smoother internal transitions raise sonority.
Table 4: Meter interaction as phonetic affordance. The table summarizes corpus-supported genre and period associations together with recurrent sonic possibilities, not fixed meter essences.
Meter Genre / historical association Recurrent phonetic affordance Illustrative poets
Mutaqarib Epic masnavi, heroic narration, and later didactic-public reuse; especially prominent in early classical epic memory Repeated cadential closure and compact consonantal sequencing support firmness and cluster-backed assertion more readily than in lyric-dominant meters. Ferdowsi, Saadi, Bolandeghbal
Hazaj Romance masnavi, chantable lyric extension, and later devotional or pedagogical reuse Favors smoother linear motion and higher sonority, yet can still carry sibilant incision when devotional diction sharpens the line. Fakhr Al-Din Asad Gorgani, Safi Alishah, Jahan Malek Khatun
Ramal Ghazal, mystical lyric, and reflective didactic verse; especially central from the high-classical lyric centuries onward Preserves strong contrast on the hardness-sonority axis rather than forcing a single texture; contemplative softness and pressure-rich intricacy can coexist within the same envelope. Shah Nematollah Vali, Bidel Dehlavi, Khaqani
Mujtass Compact ghazal and courtly lyric, particularly in performance-oriented lyric repertoires Encourages openness, vowel prominence, and local contour, making it hospitable to bright lyric memorability without excessive compaction. Hafez, Jahan Malek Khatun, Ashofteh Shirazi
Muzari’ Rhetorical lyric and learned late-classical, Indo-Persian, and revivalist textures Heightens trade-offs between hardness, contour, and entropy; poets can remain firmness-forward, contour-forward, or diversity-rich while sharing the same metrical frame. Khaqani, Saadi, Bidel Dehlavi, Ashofteh Shirazi
Table 4 condenses the point in literary-historical terms. The same meter can host different poets and genres, but not without consequence: each bahr offers a bounded set of rhythmic and phonotactic opportunities. Mutaqarib stabilizes firmness for epic and didactic declaration, hazaj accommodates both romance smoothness and devotional insistence, ramal allows contemplative openness as well as pressure-rich intricacy, mujtass favors bright lyric contour, and muzari’ sharpens rhetorical contrast. Meter therefore mediates phonetic style without determining it.
The within-meter evidence supports the same interpretation. Mean pairwise correlation of poet effects is 0.58 for hardness, 0.68 for sonority, and 0.31 for sibilance. Sonority is therefore the most portable dimension across metrical envelopes, hardness remains moderately portable, and sibilance is the most meter-sensitive. That pattern is historically plausible. Vocalic openness can travel across several repertoires, whereas strident contour tends to sharpen most strongly when a meter already licenses it. The full within-meter standardized-effects tables are therefore moved to Appendix B, where readers can inspect the complete coefficient matrices for hardness, sonority, and sibilance without interrupting the main historical argument.
7 Case Studies: Poet Profiles
The case studies convert the comparative field into literary argument. Ten poets belong to the balanced inferential cohort and therefore support controlled or within-meter interpretation. Three further anchors, Ferdowsi, Shahriar, and Khayyam, enter descriptively from the wider poet atlas because any long historical account of Persian poetic sound would be distorted without epic, modern lyric, and rubai as reference points. The aim is not to turn every poet into a miniature coefficient table, but to read phonetic profiles alongside genre, performance context, and literary-historical reputation. The broader fingerprint atlas for the full comparative field appears in Appendix A.
Figure 7: Case-study fingerprint atlas for thirteen focal poets and comparative anchors. Every panel uses the same metric order and the same 0–1 radial normalization derived from the full poet atlas, so differences can be compared directly across poets rather than inferred from separate scales. Ferdowsi, Saadi, and Khayyam lean toward hardness and closure; Hafez and Jahan Malek Khatun toward sonority and vowel spread; Safi Alishah and Khaqani toward sibilant contour; Bidel toward entropy and cluster pressure; and Shahriar toward modern descriptive openness.
Figure 7 provides the visual entry point for the case analyses. Because every poet is plotted on identical axes, the central trade-offs can be read before the prose narrows to individual authors. The figure also clarifies why a single hard-soft vocabulary is inadequate: Hafez and Shah Nematollah Vali are both open, but not in the same way; Ferdowsi and Saadi are both firm, but for different literary functions; Safi Alishah and Khaqani both intensify contour, but not for the same rhetorical ends.
Table 5: Case-study meter invariance summary for the balanced inferential subset only. Each cell reports standardized within-meter effects for Hardness/Sonority/Sibilance; dashes indicate poet-meter cells without sufficient mesra support. Columns correspond to the retained named meters mutaqarib, hazaj, ramal, mujtass, and muzari’. Descriptive anchors such as Ferdowsi, Shahriar, and Khayyam are discussed in the text but do not appear here because the table is restricted to supported within-meter estimates.
Poet Mutaqarib Hazaj Ramal Mujtass Muzari’
Hafez 0.03/0.17/-0.03
Khaqani 0.04/-0.00/0.00 0.07/-0.06/0.09
Saadi 0.25/-0.11/-0.01 0.10/0.03/-0.24 0.21/-0.12/-0.09
Bidel Dehlavi 0.22/-0.15/0.06 0.13/-0.06/0.02 0.15/-0.24/0.06
Safi Alishah 0.06/-0.20/0.08
Shah Nematollah Vali -0.35/0.26/-0.10
Ashofteh Shirazi 0.05/-0.08/-0.08 0.15/-0.22/0.04
Asad Gorgani -0.03/0.32/-0.23
Jahan Malek Khatun -0.01/0.29/-0.08 -0.01/0.35/-0.03 0.18/0.16/0.11
Bolandeghbal -0.06/-0.02/0.13 0.00/-0.04/-0.04
Table 5 summarizes only the ten balanced case-study poets for whom within-meter estimation is adequately supported. Ferdowsi, Shahriar, and Khayyam remain in the section as historically necessary descriptive anchors, but they are not folded into unsupported inferential claims. The distinction matters because the argument of the article depends on keeping literary centrality and evidential support in view together.
Figure 8: Within-meter comparison for the supported case-study subset. Columns correspond to the retained classical meters mutaqarib, hazaj, ramal, mujtass, and muzari’, and rows to focal poets with sufficient mesra support in at least one cell. The three aligned heatmaps report standardized within-meter effects for hardness, sonority, and sibilance; red cells indicate positive deviations, blue cells negative deviations, and gray cells unsupported poet-meter combinations. The figure makes visible Saadi’s portable hardness, Bidel’s multiaxial contrast across several meters, Shah Nematollah Vali’s strong ramal reversal toward low hardness and high sonority, and the more localized contour signatures of Hafez and Safi Alishah.
Figure 8 shifts from poet fingerprints to meter-conditioned portability. Rather than tracing poem progression, it asks which phonetic tendencies survive when poets are compared inside the same aruz frame. The result clarifies the difference between broad stylistic identity and meter-specific support: Saadi and Bidel retain structured profiles across multiple meters, Hafez concentrates in a single supported mujtass cell, Safi Alishah is sharply defined in hazaj, and Shah Nematollah Vali’s negative-hardness counterprofile is decisively a ramal effect. Gray cells should therefore be read as absent support, not as neutral style.
7.1 Safi Alishah
Historical context. Safi Alishah belongs to the late Qajar world of Ni’matullahi Sufi pedagogy, devotional exposition, and print-era circulation. His verse is less mannerist than exhortative: it aims at mystical instruction that can also be remembered and recited. Fingerprint. That expectation is only partly confirmed by the profile: hardness stays close to the corpus center, sonority is clearly depressed, and sibilance rises to one of the strongest positive levels in the case panel. The within-meter estimate in the hazaj subset repeats the same pattern (0.06/0.20/0.080.06/-0.20/0.08 for hardness/sonority/sibilance), which matters because hazaj often carries smoother romance or devotional continuity. Comparison. Safi is therefore closer to Khaqani than to Shah Nematollah Vali on the contour axis, but without Khaqani’s courtly entropy.
Interpretation. The result is a poetics of exhortative sharpness rather than block-like pressure: edge is produced by articulated stridency, not by maximal hardness. In literary terms, this suits a pedagogical Sufi voice that must remain intelligible while still sounding urgent enough for admonition and remembrance. Example. Representative devotional phrases such as tahaqqoq bar sefaat-e haqq-e vaahed gather repeated /h/, /q/, and /s/ sounds into a narrow, incisive texture.
7.2 Shah Nematollah Vali
Historical context. Shah Nematollah Vali stands at the center of Persian Sufi institutional history as the founding figure of a major mystical lineage, and his poetic reception is tied to devotional intimacy, ethical instruction, and transmissible spiritual counsel. Fingerprint. He anchors the negative-hardness pole in the controlled cohort: hardness is very low, sonority is high, vowel ratio is elevated, and the within-meter ramal estimate shows the strongest supported reversal in the study (0.35/0.26/0.10-0.35/0.26/-0.10). Because the balanced corpus places him almost entirely in ramal, the profile should be read as a distinctive exploitation of one of Persian lyric’s most flexible mystical meters rather than as meter-free softness. Comparison. He shares openness with Asad Gorgani and Jahan Malek Khatun, but differs from both by pairing that openness with higher entropy and a clearer mystical-pedagogical cadence.
Interpretation. Low hardness here is not absence of structure. It is an alternative structure in which sonorous breadth and internal variation carry intensity, fitting a devotional style that seeks persuasive continuity rather than aphoristic closure. The metrics therefore give acoustic form to the difference between mystical counsel and rhetorical declaration. Example. The poet’s characteristic lines rely on liquids, nasals, and open vowels rather than on hard consonantal blockage, which helps explain why contemplative insistence can sound fluid rather than blunt.
7.3 Bidel Dehlavi
Historical context. Bidel is a defining figure of the later Indo-Persian sabk-e hindi, a style repeatedly described as difficult, compressed, and conceptually overcharged. Later criticism treats that difficulty as imagistic and cognitive as much as verbal. Fingerprint. The phonetic profile gives that difficulty a more precise shape. Bidel is not simply “hard.” His most striking value is entropy, which is among the highest in the corpus, accompanied by cluster support, moderate positive hardness, and negative sonority. The meter-stratified estimates remain notably stable across ramal, mujtass, and muzari’. Comparison. Relative to Saadi, Bidel is less aphoristically firm and more internally variegated; relative to Khaqani, he is less specialized in sibilance and more strongly defined by diversity.
Interpretation. The style’s pressure comes from rapid alternation and symbolic heterogeneity rather than from one dominant consonantal register. This matters literarily because it shows how later Persianate difficulty can be carried by sound without reducing that difficulty to sheer harshness. The relative stability across three meters also suggests that Bidel’s density is portable rather than tied to one exceptional prosodic shell. Example. Fragments such as gar qadat kham kard… already show Bidel’s tendency to let repeated consonants and shifting sonority coexist in one tightly packed line.
7.4 Ashofteh Shirazi
Historical context. Ashofteh Shirazi belongs to the Qajar bazgasht-e adabi milieu, where poetic authority is tied to classical legibility, prosodic training, and rhetorical polish rather than to avant-garde difficulty. Fingerprint. His profile is mixed rather than extreme: hardness is modestly positive, sonority sits near the center descriptively but becomes negative under control, vowel ratio and entropy are both elevated, and the mujtass and muzari’ estimates preserve that moderated tension. Comparison. Ashofteh lies between Safi’s sharper devotional contour and Bidel’s denser complexity. He is also less sonority-led than Jahan Malek Khatun and less hardness-led than Saadi.
Interpretation. This is not a bland middle case. It is a style of managed contrast in which firmness, flow, and symbolic variety are held in calibrated balance, which suits a return-style poetics that revives classical resources without simply replicating earlier lyric smoothness. The important point is that bazgasht does not flatten the acoustic field; it reorganizes it into a more legible but still varied classical surface. Example. Lines of the form har kas ke goft … az khataa kojaast illustrate how repeated /kh/, /g/, and /t/ can sustain edge without destroying lyrical motion.
7.5 Fakhr Al-Din Asad Gorgani
Historical context. Asad Gorgani’s place in Persian literary history is secured above all by the romance tradition of Vis o Ramin, where narrative continuity and sensuous extension matter as much as rhetorical brilliance. Fingerprint. The profile is accordingly sonority-forward: hardness is below baseline, sonority and vowel ratio are elevated, sibilance is suppressed, and the within-meter hazaj estimate shows one of the clearest positive sonority effects in the case panel (0.03/0.32/0.23-0.03/0.32/-0.23). Because his surviving corpus in the balanced frame is almost entirely hazaj, the profile captures not only romance diction but a meter historically hospitable to melodic continuity. Comparison. Asad shares openness with Shah Nematollah Vali and Jahan Malek Khatun, but his lower entropy makes the texture smoother and more architectonic.
Interpretation. The result is not softness as weakness, but narrative smoothness as formal continuity. The low-entropy value is especially important: it distinguishes romance fluency from later lyric or mystical openness by showing that the line remains orderly even when it is sonority-rich. Example. Romance lines such as ke bar bidaad-e to del sakht karda-st show that moments of pressure occur, yet the larger profile remains one of vocalic extension rather than sustained compaction.
7.6 Ferdowsi
Historical context. Ferdowsi is the indispensable epic reference point of Persian literary history. The Shahnameh does not represent only one poet or one book; it establishes a model of heroic narration, dynastic memory, martial declaration, and elevated narrative steadiness that later Persian poetry continually remembers, revises, or resists. Fingerprint. In the broader atlas he occupies the hard-rhetorical quadrant: hardness is elevated, cluster support is positive, sonority is relatively restrained, and sibilance does not dominate the profile. Comparison. He therefore stands near Saadi and descriptively near Khayyam on the firmness axis, but differs from both by the scale and public amplitude of epic narration.
Interpretation. Ferdowsi’s elevated hardness reflects the epic register and heroic diction of the Shahnameh. What the metrics capture is not bluntness for its own sake, but an architecture of closure, stress, and forward thrust suited to battle scenes, royal speeches, genealogical transitions, and acts of sovereign decision. Unlike later rhetorical poets, he does not need high sibilance to sound forceful. Epic authority here is carried by consonantal steadiness and metrical drive rather than by bright strident contour.
7.7 Saadi
Historical context. Saadi is central to Persian ethical, didactic, and lyric tradition, and his reputation depends on fasahat, balance, tonal range, and aphoristic authority rather than on obscurity. Fingerprint. He is also one of the strongest hardness-forward poets in the controlled frame. Descriptive hardness is high, sibilance is below average, and within-meter estimates keep hardness positive across mutaqarib, mujtass, and muzari’. Comparison. Saadi therefore differs sharply from Safi and Khaqani, whose edge is more contour-driven, and from Hafez, whose lyric distinctiveness depends more on openness and selective sibilance.
Interpretation. Saadi’s firmness is architectural: consonantal closure and compact transitions carry authority without requiring strong strident color. Its persistence across epic, lyric, and rhetorical meters shows that this is not simply one register effect. The profile gives quantitative form to a longstanding critical intuition that Saadi’s lucidity is disciplined rather than acoustically diffuse. Example. The famous line tariqat be joz khedmat-e khalq nist embodies that structure through repeated /t/, /k/, and /kh/ sounds that reinforce lucid severity.
7.8 Khaqani
Historical context. Khaqani is the paradigmatic learned qasida poet: allusive, difficult, syntactically compressed, and rhetorically competitive. In his case, difficulty is part of poetic status, not a by-product. Fingerprint. The controlled profile confirms distinction but shifts its mechanism. Hardness is near the center, whereas sibilance and entropy are more diagnostic. The ramal and muzari’ estimates preserve mild hardness and clearer contour, especially in the latter meter. Comparison. Khaqani is nearer to Safi Alishah on the sibilance axis than to Saadi on the hardness axis, but he differs from Safi by greater lexical and phonetic complexity.
Interpretation. Courtly force here is carried less by massed obstruence than by precision, brightness, and distributional intricacy. That result helps explain why Khaqani often feels difficult through compression and brilliance rather than through simple roughness. It also clarifies why qasida rhetoric should be treated as articulated brightness as much as consonantal pressure. Example. Phrases like tahdid-e tigh mi-dahad… show how repeated /t/, /d/, /gh/, and /sh/-adjacent textures can sharpen rhetorical display without pushing the poet to the hardness extreme.
7.9 Jahan Malek Khatun
Historical context. Jahan Malek Khatun, the poet-princess of fourteenth-century lyric culture, is primarily associated with ghazal, courtly refinement, and emotionally articulate address within classical forms; her surviving divan is also unusually large for a premodern woman poet. Fingerprint. Her profile is one of the strongest sonority-vowel configurations in the case panel: high sonority, high vowel ratio, low cluster pressure, and positive within-meter sonority in hazaj, mujtass, and muzari’. Comparison. She shares openness with Hafez and Shah Nematollah Vali, but differs from Hafez by milder contour and from Shah Nematollah Vali by less extreme hardness reversal.
Interpretation. The result is lyric luminosity rather than mystical softness alone. Her lines stay open and vocally expansive while retaining enough consonantal contour to avoid diffusion, which helps situate her between courtly polish and intimate lyric release. The breadth of her meter distribution matters here because it shows that this lyric openness is not confined to a single prosodic niche. Example. A fragment such as che khosh vaqt ast vaqt-e gol be bostaan illustrates how repeated vowels and liquids create an audible sense of release.
7.10 Bolandeghbal
Historical context. Bolandeghbal belongs to a later classical-revival environment in which didactic and devotional registers remain active but are often delivered through patterned public address and Sufi-ethical exhortation. He is analytically useful precisely because he represents a later admonitory voice that is less canonically theorized than Saadi or Hafez but still metrically well supported. Fingerprint. The descriptive profile sits near the center overall, yet the within-meter evidence reveals a more pointed structure: in mutaqarib he shows mild negative hardness with positive sibilance, while ramal pulls him toward a quieter center. Comparison. He is therefore less extreme than Safi on contour and less firmness-driven than Saadi, but he still belongs near the contour-oriented side of the stylistic map.
Interpretation. The profile suggests Sufi-didactic diction shaped by repeated edge effects rather than by sustained compaction. The split between mutaqarib and ramal indicates that his audible bite is selective rather than global, which is exactly what one would expect from admonitory verse that sharpens at moments of warning, counsel, or public address. Example. Lines such as dare baagh raa baaghbaan sakht bast display the recurrent play of /b/, /gh/, /s/, and /kh/ that gives the voice its admonitory snap.
7.11 Shahriar
Historical context. Shahriar stands at the threshold of modern Persian lyric, where classical ghazal survives but does so under stronger colloquial pressure, performance modernity, and an altered public soundscape. Modern criticism repeatedly links his idiomatic diction to conscious consonantal and assonantal patterning. Fingerprint. In the broader corpus, his descriptive profile is highly distinctive: hardness is low, sonority and vowel ratio are high, entropy remains positive, and cluster pressure is notably reduced. Comparison. Shahriar resembles Jahan Malek Khatun and Shah Nematollah Vali in openness, but he moves further toward conversational spaciousness and away from classical compaction.
Interpretation. Because he falls outside the balanced inferential cohort, the profile must be read as descriptive rather than as a controlled poet effect. Even so, it strongly suggests a modernizing sonorous openness in which expressive immediacy is carried by vowel spread, reduced consonantal density, and selective patterning rather than by the heavily wrought closure of classical diction. The pattern suits a twentieth-century lyric culture in which recital, song adaptation, and conversational address reshape what counts as persuasive poetic sound. Example. The relevant sonic effect is less a single famous cluster than a general relaxation of the classical pressure profile.
7.12 Hafez
Historical context. Hafez is the canonical ghazal poet of singable ambiguity, iham, lyric memorability, and carefully wrought sound patterning. His poems are historically associated with performance as much as reading, and criticism often stresses deliberate extra-metrical sonority. Fingerprint. His profile combines high sonority, high vowel ratio, elevated sibilance, and unusually low cluster pressure. The pooled controlled model identifies a mixed lyric signature, while the single supported within-meter cell in mujtass keeps sonority positive and hardness mildly positive. Comparison. Hafez shares openness with Jahan Malek Khatun and Shah Nematollah Vali, but carries more contour than either; he shares contour with Khaqani, yet in a much lower-compaction lyric register.
Interpretation. The result is a hinge style between sonority-led lyricism and sibilant contouring. Quantitatively, that hinge helps explain why Hafez can sound both fluid in recitation and sharply memorable in local phrasing. It gives numerical support to the longstanding critical claim that Hafez’s ghazals are singable without losing epigrammatic bite, a combination that also helps explain their exceptional afterlife in musical and oral performance. Example. Phrases such as hadis-e ’eshq ze Haafez shenow show how repeated /h/, /sh/, /s/, and open vowels can make a line fluid and sharply memorable at once.
7.13 Khayyam
Historical context. Khayyam’s literary identity is inseparable from the rubai and from the long history of attributional instability that surrounds the corpus. That instability is not a marginal footnote here; it is part of what makes Khayyam a useful descriptive edge case. Fingerprint. In descriptive terms, the broader profile is firmness-oriented: hardness is above average, cluster support is positive, sonority is near the center, and entropy is slightly reduced. Comparison. Khayyam thus belongs nearer to Saadi than to Hafez or Shahriar on the openness-pressure field, but without Saadi’s broader controlled support.
Interpretation. Because the balanced inferential frame does not estimate a comparable controlled effect, this profile should not be treated as definitive authorial essence. It is better read as evidence that the Khayyamic tradition favors compact closure and epigrammatic firmness, which suits the argumentative turn structure of the rubai. At the same time, any divergence from expectation may reflect the composite history of the attributed corpus as much as a single authorial voice. Example. The relevant acoustic economy is that of brevity itself: lines that arrive quickly at closure and rely on pressure rather than expansive lyric spread.
7.14 Cross-Case Synthesis
The thirteen cases show why a serious digital-humanities treatment of poetic sound must remain both quantitative and literary-historical. Similar literary labels can be realized through different phonetic pathways. Mystical or lyric poetry can be sonority-led, as in Shah Nematollah Vali or Jahan Malek Khatun, but it can also be contour-led, as in Hafez or Safi Alishah. Rhetorical force can be hardness-dominant, as in Ferdowsi or Saadi, or sibilance-dominant, as in Khaqani. Complexity can arise through entropy and local contrast rather than through generalized roughness, as Bidel makes clear, while return-style classicism can preserve legibility without flattening diversity, as Ashofteh shows.
Across the panel, Persian prosody appears not as inert backdrop but as mediating institution. Mutaqarib supports firmness for Saadi and, in the longer epic tradition, forms the natural backdrop against which Ferdowsi is heard; hazaj sustains both Asad Gorgani’s romance smoothness and Safi Alishah’s devotional edge; and ramal houses both Shah Nematollah Vali’s contemplative openness and Bidel’s density. Mujtass sharpens the ghazal contour of Hafez and Jahan Malek Khatun, while muzari’ becomes a site of later rhetorical or mixed-regime contrast. The same meter can therefore host divergent phonetic strategies, but only within a constrained prosodic envelope. This is where the computational findings become most legible to literary history: style emerges as deviation, intensification, or recombination within inherited rhythmic forms.
Taken together, the case studies show that familiar literary labels such as smoothness, severity, luminosity, or difficulty are not acoustically uniform categories. Epic firmness, lyrical openness, mystical contour, and later Persianate density emerge through different phonetic routes, and the framework makes those routes comparable without flattening their historical distinctiveness.
8 Discussion
The central result of this study is that phonetic style in Persian poetry is measurable, historically situated, and formally mediated. Poet-level differences remain visible after control, but those differences are moderate in magnitude and uneven in portability across named metrical environments. This dual result rejects two simplifications at once: a formalist reduction in which meter exhausts style, and an essentialist model in which each poet carries an acoustically self-sufficient identity.
8.1 Style as Configuration Rather Than Scalar Rank
One of the main interpretive gains of the article is the shift from scalar ranking to profile configuration. Terms such as smoothness, severity, luminosity, density, or brilliance are common in Persian literary criticism, but they compress several distinct phonetic routes. A poet may sound forceful through hardness, through sibilant contour, through cluster pressure, or through a combination of these. A poet may sound open through sonority, through vowel spread, through reduced compaction, or through all three.
The fingerprints make those routes visible. Saadi’s authority is hardness-led rather than strongly sibilant. Hafez’s lyric memorability depends on openness plus contour. Shah Nematollah Vali and Jahan Malek Khatun share sonorous breadth, but not the same hardness reversal or entropy. Bidel’s difficulty arises less from blunt roughness than from internally varied pressure. Ferdowsi shows that epic authority can remain forceful without bright fricative emphasis. The value of measurement here is not to replace critical vocabulary, but to decompose it into more exact comparative terms.
8.2 Historical Formations and Performance Context
The stylistic clusters are most meaningful when understood as historical formations rather than geometric conveniences. The high-sonority zone gathers poets whose openness serves different literary institutions: Asad Gorgani’s romance narration, Jahan Malek Khatun’s courtly intimacy, Shah Nematollah Vali’s mystical counsel, and descriptively Shahriar’s modern lyric address. The high-hardness region joins epic, didactic, and aphoristic modes in poets such as Ferdowsi and Saadi, with Khayyam as a descriptive rubai edge case. A sibilant zone linking Safi Alishah, Khaqani, and Bolandeghbal shows that contour can underwrite both courtly brilliance and Sufi-didactic urgency. The high-entropy region of Bidel and Ashofteh marks later styles in which internal modulation itself becomes an audible resource.
Performance practice helps explain why these distributions matter. Persian poetry has long circulated through recitation, courtly assembly, pedagogical transmission, devotional repetition, and later print and public reading. Hardness, sonority, and sibilance are therefore not abstract numbers detached from literary use. They index how a line projects authority, memorability, chantability, or conceptual pressure in specific settings. Hafez’s openness and contour suit singable ghazal circulation; Safi Alishah’s narrowed sonority and higher sibilance suit admonitory Sufi address; Ferdowsi’s compact firmness suits sustained heroic narration; Shahriar’s vowel-rich spaciousness suits a modern lyric heard in a different public soundscape.
8.3 Prosody as Mediating Institution
The results also clarify how prosody mediates literary history. Persian meter should not be treated as a detachable wrapper around diction. A bahr is a quantitative pattern of long and short syllables, not a vague tonal mood, and that pattern creates recurring opportunities for closure, openness, contour, and diversity. The dominant named bahrs retained in this study, together with their specific sub-patterns, therefore shape phonetic opportunity before any individual poet is heard. The five-way statistical grouping used here is a post-hoc analytic convenience rather than a canonical literary category beyond those named meters themselves. This is why meter control strengthens rather than weakens interpretation: it distinguishes inherited form from stylistic work within form.
The century results reinforce the same point at a larger scale. Early corpora retain a firmer epic and courtly baseline; the thirteenth through fifteenth centuries expand sonorous lyric and mystical diction; the sixteenth and seventeenth centuries intensify entropy and mixed pressure in later Persianate experimentation; and nineteenth-century revival, devotional, and didactic voices redistribute rather than erase earlier resources. The computational evidence therefore supports a historical poetics of shifting affordances rather than abrupt sonic rupture. Meter, genre, and institutional setting jointly shape what kinds of phonetic distinctiveness become durable.
8.4 Implications for Digital Humanities
For digital humanities, the broader lesson is methodological. A strong humanities-computation article should not oppose reproducibility to interpretation. It should make representation choices explicit, control design historically intelligible, and figures readable enough for scrutiny. Quantitative work is most persuasive when it identifies which differences are formally structured, historically meaningful, and worth renewed critical debate.
In that sense, the contribution of this article is not only substantive but programmatic. It suggests a way of studying poetic sound that is cumulative rather than reductive: corpus-scale enough to reveal structure, controlled enough to support comparison, and interpretive enough to remain answerable to literary history.
9 Limitations
The strongest claims of this article are deliberately tied to a restricted inferential domain. Controlled models are estimated only on the five most frequent named classical meters and on poet-meter cells with at least 2,000 mesras, so low-support poets, rare forms, and some historically important corpora fall outside the main comparison set. Ferdowsi, Shahriar, and Khayyam illustrate this boundary directly: they remain indispensable for comparison, but they should not be absorbed into the same within-meter inferential frame as the high-support cohort. The resulting effect sizes are meaningful yet moderate, and they should be read as structured stylistic tendencies rather than as exhaustive summaries of poetic value.
The analysis is also mediated by representation and corpus history. Persian orthography omits many short vowels, so the grapheme-to-phoneme layer remains a disciplined reconstruction rather than a record of historical recitation; meter, form, and line length are controlled, but finer performative and regional conditions are not. Attribution uncertainty, genre concentration, and canon formation further shape which poets survive in volumes large enough for stable estimation. The framework is therefore best understood as a confound-aware comparative instrument that invites future genre-stratified, attribution-sensitive, and performance-aware extension.
10 Conclusion
This article has argued that phonetic texture in Persian poetry can be studied as a serious digital-humanities object when representation, confound control, and literary interpretation are kept in view together. The central empirical finding is not that poets occupy acoustically isolated worlds, but that measurable phonetic differentiation persists within shared formal institutions. Meter and form matter strongly; poet identity still matters after they are controlled. The result is a historically regulated field of variation rather than either formal determinism or authorial essence.
The study therefore contributes on four linked fronts already outlined in the introduction: a corpus-scale phonetic analysis of Persian poetic language, a confound-aware framework for stylistic comparison, a fingerprint-based account of poet-level configuration, and a reproducible digital-humanities workflow that keeps interpretive stakes visible. Empirically, the article shows that Persian poetic sound can be compared systematically without collapsing it into one index, that poet profiles are organized by trade-offs among openness, hardness, contour, entropy, and compaction within recognizable aruz repertoires, and that century-level change is better described as redistribution than rupture.
The case studies show why this matters for literary interpretation. Safi Alishah, Shah Nematollah Vali, Bidel Dehlavi, Ashofteh Shirazi, Fakhr Al-Din Asad Gorgani, Ferdowsi, Saadi, Khaqani, Jahan Malek Khatun, Bolandeghbal, Hafez, Shahriar, and Khayyam instantiate different phonetic pathways through epic, lyric, rhetorical, narrative, devotional, revival, and modern traditions. The contrast among Ferdowsi, Hafez, and Shahriar is especially revealing: heroic firmness, classical lyric contour, and modern vocalic openness occupy different but historically connected positions in the same phonetic field.
For digital humanities, the broader implication is methodological. A rigorous computational poetics of Persian literature should not oppose measurement to interpretation. It should use measurement to discipline interpretation, and interpretation to keep measurement historically meaningful. Under that standard, phonetic analysis becomes more than a technical supplement to stylometry: it becomes a way of hearing the corpus historically.
The restriction to five meters is therefore substantive rather than merely procedural. The retained bahrsmutaqarib, hazaj, ramal, mujtass, and muzari’—are historically documented classical meters, not ad hoc metrical families, and together they account for 77.6% of mesras with usable meter/form metadata and the entirety of the final controlled cohort. That coverage makes them the appropriate basis for corpus-wide comparison in this article: they dominate the metrically annotated archive, span epic, romance, lyric, mystical, rhetorical, and revivalist uses, and support generalization without collapsing rarer meters into spurious aggregate categories.
Acknowledgements
We are deeply grateful to the Ganjoor platform for providing open access to a rich corpus of Persian poetry, which made this computational and historical analysis possible. Their dedication to preserving and sharing these texts has been invaluable to scholars and enthusiasts alike.
This work is respectfully dedicated to the enduring voices of Ferdowsi, Saadi, Hafez, Khayyam, Khaqani, Shahriar, Safi Bolandeghbal, Jahan Malek Khatun, and other classical masters, whose eloquence, rhythm, and insight continue to illuminate the depths of Persian literary tradition. Their poetry remains a testament to creativity, perseverance, and the power of language across centuries.
Finally, we dedicate this work to the resilient people of Iran, who have endured historical and contemporary challenges with courage and dignity. We honor those who have suffered or lost their lives in the pursuit of freedom and justice, and we hope this study contributes in some small way to the celebration and understanding of their cultural and literary heritage.
Appendix A Poet Fingerprint Atlas
This appendix presents the broader fingerprint atlas used to situate the main case studies against the wider comparative field. Each sheet preserves a shared radial scale across hardness, sonority, sibilance, vowel ratio, entropy, and consonant-cluster ratio, while grouping between eight and twelve poets per page to keep the atlas readable at journal scale. Read the pages comparatively: the point is not to isolate single poets, but to see how epic, lyric, courtly, mystical, revival, and modern corpora redistribute the same phonetic resources.
Poet Profile Index
Poet Mesras
Attar 109514
Ferdowsi 99220
Saeb Tabrizi 96672
Iranshan 41650
Khwaju Kermani 41366
Nezami 33166
Elhami Kermanshahi 32174
Bidel Dehlavi 31238
Jami 29368
Amir Moezzi 24346
Rumi 21628
Qaaani Shirazi 20230
Anonymous Faramarznama 18616
Fakhr Al-Din Asad Gorgani 18004
Amir Khusrow Dehlavi 17738
Asadi Tusi 17678
Nezari Qohestani 17438
Salman Savaji 15628
Saadi 14958
Ahli Shirazi 14886
Adib Al-Mamalek Farahani 13452
Malek Al-Shoara Bahar 13352
Kamal Al-Din Ismail 12708
Ibn Yamin 12344
Salim Tehrani 11616
Fayyaz Lahiji 11094
Mohtasham Kashani 10924
Vahshi Bafqi 10798
Safi Alishah 10716
Masud Saad Salman 10320
Jahan Malek Khatun 10200
Othman Mokhtari 10176
Juyai Tabrizi 10048
Molla Masih Panipati 9988
Suzani Samarqandi 9814
Naziri Nishaburi 9764
Saif Farghani 8938
Qatran Tabrizi 8804
Khaqani 8666
Sanai 8358
Mohammad Kusaj 8150
Anvari 8016
Kalim Kashani 7672
Seyda Nasafi 7546
Rashid Al-Din Vatvat 7468
Feyz Kashani 7194
Salimi Jaruni 6886
Hazin Lahiji 6824
Bolandeghbal 6558
Adib Saber 6558
Azar Bigdeli 5844
Afsar Al-Moluk Ameli 5630
Ashofteh Shirazi 5528
Samet Boroujerdi 5054
Qodsi Mashhadi 4692
Asir Shahrestani 4504
Ayyuqi 4476
Athir Akhsikati 4376
Fuzuli Baghdadi 4342
Saghir Isfahani 4314
Mirza Aqa Khan Kermani 4248
Vaez Qazvini 4108
Farrokhi Sistani 3822
Naser Bukharaei 3688
Majd Hamgar 3500
Sheikh Mahmoud Shabestari 3368
Orfi Shirazi 3300
Jamal Al-Din Abd Al-Razzaq Isfahani 3154
Onsori Balkhi 3102
Iqbal Lahori 2996
Baba Faghani Shirazi 2974
Ohadi Maraghei 2834
Toghral Ahrari 2762
Zahir Faryabi 2754
Fayez 2552
Helali Jaghatai 2506
Hafez 2416
Safa Isfahani 2324
Qaem Maqam Farahani 2260
Torki Shirazi 2164
Shah Nematollah Vali 2120
Figure 9: Appendix atlas page 1, covering Attar through Qaaani Shirazi. The fixed radial scale allows direct comparison of openness, firmness, contour, entropy, and cluster pressure across poets from different periods and literary registers.
Figure 10: Appendix atlas page 2, covering Anonymous Faramarznama through Ibn Yamin. The fixed radial scale allows direct comparison of openness, firmness, contour, entropy, and cluster pressure across poets from different periods and literary registers.
Figure 11: Appendix atlas page 3, covering Salim Tehrani through Naziri Nishaburi. The fixed radial scale allows direct comparison of openness, firmness, contour, entropy, and cluster pressure across poets from different periods and literary registers.
Figure 12: Appendix atlas page 4, covering Saif Farghani through Hazin Lahiji. The fixed radial scale allows direct comparison of openness, firmness, contour, entropy, and cluster pressure across poets from different periods and literary registers.
Figure 13: Appendix atlas page 5, covering Bolandeghbal through Saghir Isfahani. The fixed radial scale allows direct comparison of openness, firmness, contour, entropy, and cluster pressure across poets from different periods and literary registers.
Figure 14: Appendix atlas page 6, covering Mirza Aqa Khan Kermani through Ohadi Maraghei. The fixed radial scale allows direct comparison of openness, firmness, contour, entropy, and cluster pressure across poets from different periods and literary registers.
Figure 15: Appendix atlas page 7, covering Toghral Ahrari through Shah Nematollah Vali. The fixed radial scale allows direct comparison of openness, firmness, contour, entropy, and cluster pressure across poets from different periods and literary registers.
Appendix B Meter-Stratified Coefficient Tables
These tables report within-meter standardized poet effects for the three focal metrics. They are cited in the main text whenever meter portability is discussed and are included here so that readers can inspect the evidential base behind the summary statements on hardness, sonority, and sibilance across mutaqarib, hazaj, ramal, mujtass, and muzari’.
Within-Meter Standardized Effects: Hardness
Columns correspond to mutaqarib, hazaj, ramal, mujtass, and muzari’.
Poet Mutaqarib Hazaj Ramal Mujtass Muzari’
Adib Al-Mamalek Farahani 0.00 0.00 0.00 0.00
Adib Saber -0.15 0.00
Afsar Al-Moluk Ameli -0.06
Ahli Shirazi -0.05 0.09 0.10 0.15
Amir Khusrow Dehlavi 0.15 0.11 0.08 0.12
Amir Moezzi 0.08 0.04 0.18
Anonymous Faramarznama 0.01
Anvari 0.18 0.11 0.17
Asadi Tusi 0.20
Ashofteh Shirazi 0.05 0.15
Asir Shahrestani 0.04 0.03
Athir Akhsikati 0.07 0.15
Attar 0.02
Ayyuqi -0.02
Azar Bigdeli -0.03
Baba Faghani Shirazi 0.13
Bidel Dehlavi 0.22 0.13 0.15
Bolandeghbal -0.06 0.00
Elhami Kermanshahi -0.04
Fakhr Al-Din Asad Gorgani -0.03
Farrokhi Sistani 0.04
Fayez -0.09
Fayyaz Lahiji 0.09 0.03 0.09
Ferdowsi 0.10
Feyz Kashani -0.05 -0.06 -0.01
Fuzuli Baghdadi 0.01 -0.02
Hafez 0.03
Hazin Lahiji 0.02 -0.02
Helali Jaghatai 0.09
Ibn Yamin 0.01 0.03 0.08
Imadaddin Nasimi -0.11
Iqbal Lahori -0.12
Iranshan 0.05
Jahan Malek Khatun -0.01 -0.01 0.18
Jamal Al-Din Abd Al-Razzaq Isfahani 0.16
Jami 0.06 0.05 0.04 -0.02 0.15
Juyai Tabrizi -0.00 -0.14 -0.00
Kalim Kashani 0.14 0.17 0.02
Kamal Al-Din Ismail 0.15 0.28
Khaqani 0.04 0.07
Khwaju Kermani 0.05 0.03 -0.01 0.16
Majd Hamgar 0.02
Malek Al-Shoara Bahar 0.13 0.12 0.09 0.14
Masud Saad Salman 0.00 0.14
Mirza Aqa Khan Kermani -0.04
Mohammad Kusaj -0.03
Mohtasham Kashani 0.02 -0.05 0.05
Molla Masih Panipati 0.10
Naser Bukharaei -0.08
Naziri Nishaburi 0.16 -0.01 0.05
Nezami 0.11 0.19
Nezari Qohestani 0.17 0.06 0.02 0.14
Ohadi Maraghei 0.10
Onsori Balkhi -0.03
Orfi Shirazi -0.16
Othman Mokhtari 0.02
Qaaani Shirazi 0.07 0.03 0.06
Qaem Maqam Farahani 0.03
Qatran Tabrizi -0.05 -0.03
Qodsi Mashhadi 0.15
Rashid Al-Din Vatvat -0.07 0.02
Rumi -0.08 -0.15 -0.01 -0.01
Saadi 0.25 0.10 0.21
Saeb Tabrizi 0.09 -0.00 0.12
Safa Isfahani 0.13
Safi Alishah 0.06
Saghir Isfahani -0.07 0.08
Saif Farghani 0.00 0.09 0.18
Salim Tehrani 0.01 0.07 0.02
Salimi Jaruni 0.10
Salman Savaji -0.02 0.06 0.05 0.02 0.07
Samet Boroujerdi -0.10 -0.05
Sanai -0.02
Sayyid Hasan Ghaznavi 0.10
Seyda Nasafi 0.03 -0.02
Shah Nematollah Vali -0.35
Sheikh Mahmoud Shabestari 0.06
Suzani Samarqandi -0.00 -0.05 0.03
Toghral Ahrari -0.13
Torki Shirazi 0.04
Vaez Qazvini 0.11
Vahshi Bafqi 0.06 0.15
Zahir Faryabi 0.01
Within-Meter Standardized Effects: Sonority
Columns correspond to mutaqarib, hazaj, ramal, mujtass, and muzari’.
Poet Mutaqarib Hazaj Ramal Mujtass Muzari’
Adib Al-Mamalek Farahani 0.00 0.00 0.00 0.00
Adib Saber 0.20 0.00
Afsar Al-Moluk Ameli -0.06
Ahli Shirazi -0.02 -0.01 0.02 -0.18
Amir Khusrow Dehlavi 0.11 0.15 0.11 -0.06
Amir Moezzi 0.06 0.09 -0.11
Anonymous Faramarznama 0.03
Anvari 0.06 0.15 -0.01
Asadi Tusi -0.08
Ashofteh Shirazi -0.08 -0.22
Asir Shahrestani 0.02 -0.14
Athir Akhsikati 0.15 -0.12
Attar 0.12
Ayyuqi 0.02
Azar Bigdeli -0.07
Baba Faghani Shirazi 0.00
Bidel Dehlavi -0.15 -0.06 -0.24
Bolandeghbal -0.02 -0.04
Elhami Kermanshahi 0.01
Fakhr Al-Din Asad Gorgani 0.32
Farrokhi Sistani 0.12
Fayez 0.16
Fayyaz Lahiji 0.11 0.12 -0.04
Ferdowsi 0.02
Feyz Kashani 0.08 0.05 -0.09
Fuzuli Baghdadi 0.21 0.23
Hafez 0.17
Hazin Lahiji 0.06 -0.05
Helali Jaghatai 0.08
Ibn Yamin 0.07 0.05 -0.10
Imadaddin Nasimi 0.17
Iqbal Lahori 0.21
Iranshan -0.03
Jahan Malek Khatun 0.29 0.35 0.16
Jamal Al-Din Abd Al-Razzaq Isfahani -0.02
Jami 0.05 0.15 0.16 0.22 -0.06
Juyai Tabrizi 0.04 0.12 -0.10
Kalim Kashani -0.11 0.00 0.05
Kamal Al-Din Ismail 0.04 -0.13
Khaqani -0.00 -0.06
Khwaju Kermani 0.05 0.20 0.21 -0.03
Majd Hamgar 0.15
Malek Al-Shoara Bahar -0.06 -0.02 0.01 -0.10
Masud Saad Salman 0.05 -0.01
Mirza Aqa Khan Kermani -0.06
Mohammad Kusaj 0.16
Mohtasham Kashani 0.04 0.10 -0.18
Molla Masih Panipati 0.07
Naser Bukharaei 0.21
Naziri Nishaburi 0.05 0.17 -0.04
Nezami 0.05 0.04
Nezari Qohestani -0.05 0.12 0.18 -0.03
Ohadi Maraghei -0.06
Onsori Balkhi 0.03
Orfi Shirazi 0.04
Othman Mokhtari -0.10
Qaaani Shirazi 0.08 0.16 -0.00
Qaem Maqam Farahani -0.04
Qatran Tabrizi 0.08 0.03
Qodsi Mashhadi 0.01
Rashid Al-Din Vatvat 0.24 0.06
Rumi 0.24 0.19 0.16 -0.07
Saadi -0.11 0.03 -0.12
Saeb Tabrizi 0.04 0.09 -0.19
Safa Isfahani 0.09
Safi Alishah -0.20
Saghir Isfahani 0.08 -0.11
Saif Farghani 0.14 0.10 -0.15
Salim Tehrani 0.12 0.09 0.15
Salimi Jaruni 0.00
Salman Savaji 0.09 0.10 0.17 0.14 -0.05
Samet Boroujerdi 0.17 0.12
Sanai 0.22
Sayyid Hasan Ghaznavi -0.07
Seyda Nasafi 0.07 0.06
Shah Nematollah Vali 0.26
Sheikh Mahmoud Shabestari -0.02
Suzani Samarqandi 0.11 0.18 -0.02
Toghral Ahrari 0.00
Torki Shirazi -0.07
Vaez Qazvini -0.02
Vahshi Bafqi 0.08 -0.00
Zahir Faryabi 0.14
Within-Meter Standardized Effects: Sibilance
Columns correspond to mutaqarib, hazaj, ramal, mujtass, and muzari’.
Poet Mutaqarib Hazaj Ramal Mujtass Muzari’
Adib Al-Mamalek Farahani 0.00 0.00 0.00 0.00
Adib Saber 0.13 0.00
Afsar Al-Moluk Ameli -0.06
Ahli Shirazi 0.01 -0.04 -0.10 -0.05
Amir Khusrow Dehlavi -0.18 -0.15 -0.13 -0.08
Amir Moezzi -0.08 -0.01 0.04
Anonymous Faramarznama 0.03
Anvari -0.01 -0.07 0.03
Asadi Tusi 0.17
Ashofteh Shirazi -0.08 0.04
Asir Shahrestani -0.06 0.01
Athir Akhsikati -0.03 0.01
Attar -0.21
Ayyuqi -0.03
Azar Bigdeli 0.21
Baba Faghani Shirazi -0.04
Bidel Dehlavi 0.06 0.02 0.06
Bolandeghbal 0.13 -0.04
Elhami Kermanshahi -0.02
Fakhr Al-Din Asad Gorgani -0.23
Farrokhi Sistani -0.13
Fayez 0.05
Fayyaz Lahiji -0.06 -0.08 -0.04
Ferdowsi 0.01
Feyz Kashani -0.05 0.04 0.07
Fuzuli Baghdadi -0.06 -0.07
Hafez -0.03
Hazin Lahiji 0.13 -0.09
Helali Jaghatai -0.07
Ibn Yamin 0.01 -0.02 0.02
Imadaddin Nasimi 0.09
Iqbal Lahori -0.16
Iranshan -0.00
Jahan Malek Khatun -0.08 -0.03 0.11
Jamal Al-Din Abd Al-Razzaq Isfahani -0.06
Jami 0.15 -0.06 0.01 0.03 0.12
Juyai Tabrizi -0.02 -0.09 0.03
Kalim Kashani -0.04 -0.09 -0.19
Kamal Al-Din Ismail -0.08 -0.01
Khaqani 0.00 0.09
Khwaju Kermani 0.03 0.01 -0.07 0.03
Majd Hamgar -0.05
Malek Al-Shoara Bahar 0.11 -0.01 -0.10 -0.00
Masud Saad Salman -0.09 -0.11
Mirza Aqa Khan Kermani 0.06
Mohammad Kusaj 0.01
Mohtasham Kashani -0.04 -0.02 -0.04
Molla Masih Panipati -0.11
Naser Bukharaei -0.11
Naziri Nishaburi -0.00 -0.09 -0.03
Nezami 0.03 -0.20
Nezari Qohestani 0.09 -0.16 -0.06 0.06
Ohadi Maraghei -0.01
Onsori Balkhi -0.05
Orfi Shirazi -0.11
Othman Mokhtari -0.00
Qaaani Shirazi -0.03 -0.07 0.05
Qaem Maqam Farahani -0.08
Qatran Tabrizi -0.14 -0.15
Qodsi Mashhadi 0.26
Rashid Al-Din Vatvat -0.01 0.01
Rumi -0.11 -0.03 -0.01 0.09
Saadi -0.01 -0.24 -0.09
Saeb Tabrizi -0.01 -0.09 0.03
Safa Isfahani -0.06
Safi Alishah 0.08
Saghir Isfahani -0.07 -0.00
Saif Farghani -0.08 -0.10 -0.05
Salim Tehrani 0.03 -0.05 0.01
Salimi Jaruni -0.16
Salman Savaji 0.12 -0.13 -0.04 -0.09 -0.02
Samet Boroujerdi -0.03 -0.13
Sanai -0.10
Sayyid Hasan Ghaznavi -0.08
Seyda Nasafi -0.15 -0.13
Shah Nematollah Vali -0.10
Sheikh Mahmoud Shabestari 0.01
Suzani Samarqandi -0.03 -0.03 -0.03
Toghral Ahrari -0.05
Torki Shirazi -0.04
Vaez Qazvini 0.08
Vahshi Bafqi -0.10 -0.06
Zahir Faryabi -0.12